### v1

# Self-supervised pre-training with diffusion model for few-shot landmark detection in x-ray images

Abstract
In the last few years, deep neural networks have been extensively applied in the medical domain for different tasks, ranging from image classification and segmentation to landmark detection. However, the application of these technologies in the medical domain is often hindered by data scarcity, both in terms of available annotations and images. This study introduces a new self-supervised pre-training protocol based on diffusion models for landmark detection in x-ray images. Our results show that the proposed self-supervised framework can provide accurate landmark detection with a minimal number of available annotated training images (up to $50$), outperforming ImageNet supervised pre-training and state-of-the-art self-supervised pre-trainings for three popular x-ray benchmark datasets. To our knowledge, this is the first exploration of diffusion models for self-supervised learning in landmark detection, which may offer a valuable pre-training approach in few-shot regimes, for mitigating data scarcity.
1 Introduction
Landmark detection is the task of identifying anatomical keypoints of interest in images [1]. Detecting key points can directly assist in angle measurements [2], skeletal measurements [3], and surgical planning [4].
In recent years, deep neural networks have become extensively utilized for the landmark detection task [5, 6], and many fully supervised methods have been proposed [7, 8, 9, 10, 11, 12, 13, 14, 15]. However, the limited availability of annotated data and the reliance on medical experts for annotations present challenges in applying deep learning to this domain. Specifically, annotating landmarks is labor-intensive, costly, and demands expert precision [8, 7, 9], particularly for critical tasks such as surgical treatment planning. Consequently, real-world datasets typically contain very few annotated images, requiring the design of label-efficient training solutions [16].
A transfer learning framework, typically fine-tuning ImageNet pre-trained models on medical downstream tasks, has been a common strategy to tackle this challenge [17, 18, 19]. Whether specific in-domain pre-training may benefit the model performance on the downstream task is still debated, and recent works provide controversial results depending on the specific task and domain [20, 9]. Nevertheless, involving in-domain data may be an effective alternative to ImageNet pre-training, especially in a self-supervised fashion to overcome the limited availability of annotations.
Recent advances in self-supervised learning (SSL) methods such as MoCoV3 [21], SimCLRV2 [22], and DINO [23] have shown significant potential. These approaches enable models to learn robust representations from unlabeled data, thereby reducing the dependency on large labeled training datasets. In this context, our main contribution is the proposal of a novel few-shot self-supervised pre-training approach, based on denoising diffusion probabilistic models (DDPMs), specifically focusing on landmark detection for x-ray images, widely used for anatomical assessment and diagnosis, in a realistic and typical scenario where the number of available annotations and images is minimal (up to $50$).
To our knowledge, this work is the first to employ diffusion models for self-supervision in landmark detection. We compare the proposed method with state-of-the-art alternatives including ImageNet supervised pre-training, MoCoV3, SimCLRV2, and DINO self-supervised pre-training, in terms of performance and robustness at the varying of labeled training available samples. Our results show that our method outperforms ImageNet supervised pre-training and the other self-supervised methods for the landmark detection task. Furthermore, we perform a set of experiments to provide insights on whether the usage of a different in-domain pre-training dataset may benefit the downstream task, in a few shot regimes.
The remainder of the paper is organized as follows: Section 2 reviews related studies, particularly focusing on self-supervised pre-training in the medical domain and landmark detection. Section 3 describes our proposed approach. Section 4 gives an overview of the benchmark datasets employed in this research, details the implementation specifics, and explains the evaluation metrics. Lastly, Section 5 showcases our experimental results and insights, drawing some conclusions in Section 6.
2 Related Works
The problem of few-shot learning is particularly relevant in the medical application domain, as obtaining both images and high-quality, unbiased annotations is costly and time-consuming. Consequently, several works leverage self-supervised pre-training for label efficiency [24, 25, 26, 27].
Zhu et al.[27] propose UOD (Universal One-shot Detection), a domain-adaptive framework for one-shot anatomical landmark detection across multiple medical imaging domains. UOD uses contrastive learning and a domain-adaptive transformer, while reducing annotation burden. Rousseau et al. [24] introduce a pre-training method using diffusion models for dental radiography segmentation. Baranchuk et al. [25] use diffusion models to capture high-level semantic information for semantic segmentation. Brempong et al. [26] propose Decoder Denoising Pretraining (DDeP) to enhance label efficiency in semantic segmentation models by emphasizing the importance of pre-training both encoder and decoder components with limited labeled examples.
In the specific context of landmark detection, the existing works have mainly tackled the data scarcity issue by leveraging transfer learning techniques [28, 8, 9]. Tiulpin et al. [28] use transfer learning from low-budget annotations for knee x-ray landmark localization in osteoarthritis stages. Di Via et al. [9] study whether in-domain data provide benefits over ImageNet pre-training for landmark detection in x-ray images. Zhu et al. [8] propose GU2Net, a universal model for anatomical landmark detection in medical images, addressing limitations in specialized, dataset-dependent methods. Unlike the previously cited works, in this paper, we examine the impact of self-supervised pre-training with diffusion models on landmark detection for x-ray images, focusing on a realistic and typical scenario where the number of available annotations and images is limited. To our knowledge, this is the first application of diffusion models for SSL in this specific task.
Figure 1: Schematic representation of our DDPM self-supervised landmark detection pipeline.
3 Approach
Fig. 1 illustrates our two-step methodology: initially, a DDPM is pre-trained on the training set, in a self-supervised fashion (with no annotations). Subsequently, this model is fine-tuned on the labeled training set for the downstream task, represented by the landmark detection. Further details on these steps are discussed in the following section.
Pre-training using the DDPM: Denoising Diffusion Probabilistic Models [29] (DDPMs) have emerged as a powerful class of generative models, excelling in the synthesis of high-fidelity samples in diverse fields, including medical imaging [30, 31]. Their training methodology involves a forward diffusion process that gradually transforms data into noise and a reverse process that learns to reverse this transformation, thereby generating new data samples from noise.
In this work, we leverage a DDPM Unet architecture as a predictive model for noise. During pre-training, the DDPM Unet receives batches of images alongside their respective diffusion time steps, and it is tasked with predicting the perturbation in the data. Real-world landmark detection datasets are typically small-scale, thus, we expect our DPPM Unet to be pre-trained with small-scale unannotated datasets, while still being able to provide rich and general features for the downstream task. This allows us to keep computational resources and time under a budget, while providing an effective few-shot learning approach for landmark detection.
Fine-tuning on the downstream task: The pre-trained Unet model integrated into the DDPM undergoes fine-tuning for landmark detection in a few-shot regime, thus using a few available annotated data for the downstream task. This step requires only modifying the final classification layer to align with the number of landmarks to be predicted. Although the first stage adopts a self-supervised strategy (with no annotations), the second phase employs a supervised approach by providing the model with ground-truth heat maps for each landmark to be predicted.
4 Experimental Setup
In this section, we describe the datasets used, the metrics employed, and the details of the implementation of our DDPM architectures and training procedures for the different pre-training methods discussed above. Afterwards, we show the results that have been achieved by evaluating the approach presented.
4.1 Datasets Overview
This section outlines the datasets employed in our study, including Chest x-rays, Cephalometric x-rays, and Hand x-rays. We describe the dataset composition, annotation protocols, and the division into training, validation, and testing sets to ensure clarity and reproducibility in our experiments.
Chest dataset: This dataset comprises $279$ Chest x-ray images sourced from the China set, as documented in [32], and is accessible through the U.S. National Library of Medicine’s public repository. The original image dimensions are approximately $3000$x3000300030003000 pixels. Following the procedures provided in [8, 9], we split the dataset with the first $195$ images for training, the subsequent $34$ for validation and the final $50$ for testing. We evaluate our model’s performance by pixel distance using $6$ manually annotated landmarks per image as all the PNG dataset has no information on physical spacing.
Cephalometric dataset: The Cephalometric dataset [33] is an open-source resource containing $400$ cephalometric x-ray images of individuals aged between $7$ and $76$ years. Each image has a resolution of $0.1$mmx0.10.10.10.1mm and dimensions of $2400$x1935193519351935 pixels. In our research, the first $150$ images are used for training purposes, while the remaining $250$ images are set aside for testing, following the methodology in [8]. To ensure accurate anatomical landmark annotations, two orthodontists identified $19$ landmarks on each image, and the mean coordinates were calculated to address inter-observer variability.
Hand dataset: This open dataset entails $909$ Hand x-ray images, each at an average size of $1563$x2169216921692169 pixels. Consistent with the approaches taken by [8, 9], we allocate the first $550$ images for training, the subsequent $59$ for validation, and the final $300$ images for testing. To enable a direct comparison with other approaches in terms of physical distances we take the estimate made by Payer et al.[34] that the width of the wrist is $50$mm. Then, physical distance is computed as $\frac{50}{|p-q|_{2}}$, where $p$ and $q$ are the two ends of the wrist, determined by first and fifth landmarks in the set of $37$ manually labelled by Payer et al. [34] as well.
(a) Chest landmarks: 6
(b) Cephalometric landmarks: 19
(c) Hand landmarks: 37
Figure 2: Annotated landmarks in x-ray images for Chest, Cephalometric, and Hand datasets.
4.2 Implementation Details
This section provides the technical specifics of the implementation and experimental setup. The experiments were conducted on a single NVIDIA A30 GPU with 24 GB of RAM, using Python 3.10 and PyTorch 2.1.0.
Data pre-processing: Images are resized to $256$x256256256256 pixels and normalized to a $[0-1]$ range. Post-resizing, random augmentations are applied during training, including rotations (uniformly between $-2$°and $2$°), scaling (from $-0.02$ to $0.02$), and translations (from $-0.02$ to $0.02$).
Pre-training stage: The DDPM is trained with a batch size of $4$, with gradient accumulation every $8$ batches. Training involves $500$ diffusion steps with a linear noise schedule. The training process spans $10k$ iterations, and we save the weights at iterations $4k$, $6k$, $8k$, and $10k$ for comparison during fine-tuning. We utilize the AdamW optimizer with a learning rate of $1e^{-4}$ and apply the exponential moving average to model parameters with a decay factor of $0.995$. We employ mean squared error as the loss function, consistent with the paper [29].
The DDPM Unet architecture features a multi-scale approach with channel multipliers [$1$, $2$, $4$, $8$], indicating the increase in channels at each down-sampling stage. It incorporates attention mechanisms at a resolution of $32$, uses $4$ attention heads per channel, and includes $4$ residual blocks at each resolution level.
Concerning the implementation of other SSL and ImageNet pre-training techniques, the sources used are [35] and [36].
Fine-tuning stage: The fine-tuning process is carried out using the AdamW optimizer for $200$ epochs, with an early stopping criterion. The learning rate is adjusted using a ReduceLROnPlateau scheduler based on the validation loss, starting at an initial value of $1e^{-5}$. The batch size is set at $2$, with a gradient accumulation of $8$. We use cross-entropy loss for our semantic landmark detection task to predict different classes.
4.3 Evaluation Metrics
In our research, we adopt the same evaluation metrics chosen by [8, 28, 9, 7], namely Mean Radial Error (MRE) and Success Detection Rate (SDR), which are widely used to assess the performance of landmark detection algorithms.
The MRE measures the accuracy of predicted landmarks by calculating the average Euclidean distance to the corresponding ground truth landmarks. For $n$ landmarks, the MRE is defined as:
$MRE=\frac{1}{n}\sum_{i=1}^{n}\sqrt{(x_{i}-\hat{x}_{i})^{2}+(y_{i}-\hat{y}_{i})^{2}}$
where $(x_{i},y_{i})$ represent the ground truth landmark coordinates and $(\hat{x}_{i},\hat{y}_{i})$ represent the predicted coordinates.
SDR, in contrast, evaluates robustness by indicating the percentage of predicted landmarks that fall within a certain threshold distance from the actual landmarks. For $n$ landmarks with a threshold $t$, the SDR is calculated as:
$SDR=\frac{1}{n}\sum_{i=1}^{n}I(d_{i}\leq t)$
where $d_{i}$ is the Euclidean distance between the ground truth and the predicted landmark, and $I$ is an indicator function that returns $1$ if $d_{i}\leq t$ and 00 otherwise.
5 Results and Analysis Discussion
In this section, we present the experiments conducted to evaluate our proposed methods and the corresponding results. We detail the tuning of DDPM pre-training iterations, assess the performance for landmark detection with few labeled data, and analyze the impact of different pre-training datasets on the downstream task.
5.1 Tuning the DDPM pre-training iterations
The number of training iterations during the pre-training of the DDPM is a fundamental hyperparameter of our proposed approach. To evaluate the robustness and select the number of training iterations, we perform an experiment running three times the entire pipeline, considering $4k$, $6k$, $8k$, and $10k$ training iterations for the DDPM pre-training step.
Figure 3 visualizes the landmark detection results obtained in the validation sets for the Chest, Cephalometric, and Hand datasets, in terms of MRE. Our findings suggest that the pre-training is quite robust, except for the $4k$ iterations where the model is likely still learning crucial data features (as shown by the high standard deviation). Excessive training iterations beyond $8k$ may result in overfitting, and a slightly degrading performance on the downstream task. Considering these results, we decided to select the optimal number of training iterations for DDPM pre-training for subsequent experiments, utilizing $6k$ iterations for the Chest dataset, $8k$ iterations for the Cephalometric dataset, and $8k$ iterations for the Hand dataset.
Figure 3: Landmark detection performance on Chest, Cephalometric, and Hand validation sets across DDPM pre-training iterations.
5.2 Downstream task performance evaluation
In this section, we assess the effectiveness of our DDPM self-supervised pre-training method, benchmarking it with supervised ImageNet pre-training, and self-supervised state-of-the-art methods including MoCoV3, SimCLRV2, and DINO, across different numbers of labeled training samples (up to $50$) in the Chest, the Cephalometric, and the Hand datasets. Tables 1, 2, and 3 report the average test results for three independent runs on the three datasets (a bar chart version of the tables is provided in the Supplementary Material). With the exception of our approach, which implements a DDPM as described in the original paper [29], all reported results utilize DenseNet161 as the encoder backbone. We opt for this backbone due to its performance in the landmark detection task across all the datasets tested, based on a hold-out validation comparison with VGG19 and ResNeXt50_32x4D. Our approach outperforms both ImageNet and the explored alternative SSL approaches for all the tested datasets and quantities of training images in the downstream task. These results are qualitatively supported by Figure 4, providing a visual representation (i.e., the prediction heatmaps) of the methods’ performance with $10labels$, for each of the datasets included in our work.
Performance gains are particularly pronounced in low-data regimes, crucial for medical imaging applications where labeled data is often scarce.
In the Chest dataset (Table 1), with just one labeled sample, our method achieves an MRE of $14.99$ px compared to $143.67$ px for ImageNet and $64.69$ px for the best alternative (DINO). This represents a $89.6$% reduction in MRE versus ImageNet and a $76.8$% reduction versus DINO. The SDR at $9$px shows even more improvements, with our method reaching $64.63$%, far surpassing ImageNet (00%) and DINO ($23.44$%).
Similar trends are observed in the Cephalometric dataset (Table 2). Using one labeled sample, our method attains an MRE of $6.87$ mm versus $48.69$ mm for ImageNet and $24.39$ mm for the best alternative (MoCoV3) - reductions of $85.9$% and $71.8$%, respectively. The SDR at the restrictive 2mm further highlights our method’s superiority, achieving $49.43$% compared to ImageNet’s $0.12$% and MoCoV3’s $18.36$%.
The Hand dataset (Table 3) results reinforce these findings. With one labeled sample, our approach yields an MRE of $28.75$ mm versus $79.32$ mm for ImageNet and $97.31$ mm for the best alternative (DINO) - reductions of $63.8$% and $70.5$%, respectively. The SDR at 2mm shows our method at $27.87$%, significantly outperforming both ImageNet ($0.04$%) and DINO ($0.47$%).
As the number of labeled samples increases beyond $10$, the performance gap narrows, though our approach still maintains an edge. For instance, in the Cephalometric dataset with $25$ labeled samples, our method’s MRE of $1.62$ mm is closer to, but still better than, ImageNet’s $4.18$ mm and SimCLRV2’s $2.19$ mm. This trend is consistent across all datasets, indicating that while our method provides substantial benefits in extremely low-data scenarios, for which it is conceived and proposed, it continues to offer improvements even as more labeled data becomes available.
The obtained results highlight the efficacy of our DDPM self-supervised pre-training approach for landmark detection in few-shot learning scenarios, prevalent in medical imaging contexts where available high-quality annotated images are typically limited.
Finally, to frame our results in the state-of-the-art, we can consider that the current best-performing method [9] for the landmark detection task of the whole Chest, Cephalometric, and Hand datasets, consists of fine-tuning an ImageNet pre-trained model on the target datasets.
Table 1: Comparing landmark detection test performance on the Chest dataset through our DDPM self-supervised pre-training and ImageNet, MoCoV3, SimCLRV2, DINO pre-training across varying numbers of labeled training images, highlighting best results in bold.
Number of Labeled Training Samples Pre-trained Models Chest
MRE $\downarrow$ SDR(%) $\uparrow$
(px) 3px 6px 9px
$1$ ImageNet $143.67\pm 0.27$ $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00 $00.00\pm 0$ 0.00
MoCoV3 $106.74\pm 14.95$ $2.98\pm 0.51$ $12.33\pm 2.26$ $19.92\pm 3.47$
SimCLRV2 $118.97\pm 1.84$ $1.90\pm 0.19$ $6.09\pm 0.58$ $11.39\pm 0.58$
DINO $64.69\pm 0.91$ $3.66\pm 0.33$ $15.72\pm 0.38$ $23.44\pm 0.38$
Our method $\mathbf{14.99\pm 0.00}$ $\mathbf{19.92\pm 0.00}$ $\mathbf{46.34\pm 0.00}$ $\mathbf{64.63\pm 0.00}$
$5$ ImageNet $72.81\pm 2.71$ $2.44\pm 0.33$ $8.54\pm 1.00$ $14.09\pm 1.38$
MoCoV3 $18.35\pm 0.18$ $14.63\pm 1.00$ $41.60\pm 0.69$ $66.53\pm 0.69$
SimCLRV2 $30.09\pm 0.35$ $15.72\pm 0.69$ $48.37\pm 1.45$ $67.48\pm 0.33$
DINO $24.48\pm 1.73$ $11.25\pm 1.16$ $32.65\pm 0.51$ $48.78\pm 1.00$
Our method $\mathbf{6.04\pm 0.00}$ $\mathbf{32.52\pm 0.00}$ $\mathbf{65.04\pm 0.00}$ $\mathbf{86.99\pm 0.00}$
$10$ ImageNet $35.83\pm 1.57$ $7.11\pm 0.84$ $22.56\pm 1.06$ $37.30\pm 1.39$
MoCoV3 $13.52\pm 0.67$ $18.49\pm 0.21$ $44.72\pm 1.21$ $67.28\pm 1.43$
SimCLRV2 $28.50\pm 0.00$ $17.89\pm 0.00$ $45.93\pm 0.00$ $64.63\pm 0.00$
DINO $14.96\pm 0.03$ $16.26\pm 0.41$ $39.64\pm 0.21$ $61.99\pm 0.61$
Our method $\mathbf{5.33\pm 0.00}$ $\mathbf{33.74\pm 0.00}$ $\mathbf{67.89\pm 0.00}$ $\mathbf{86.59\pm 0.00}$
$25$ ImageNet $11.38\pm 1.11$ $20.83\pm 2.00$ $49.29\pm 1.13$ $70.33\pm 0.29$
MoCoV3 $6.15\pm 0.00$ $30.89\pm 0.00$ $66.67\pm 0.00$ $85.37\pm 0.00$
SimCLRV2 $11.46\pm 0.00$ $33.74\pm 0.00$ $69.11\pm 0.00$ $84.15\pm 0.00$
DINO $10.32\pm 0.35$ $24.59\pm 1.02$ $55.69\pm 0.81$ $75.81\pm 0.20$
Our method $\mathbf{4.62\pm 0.00}$ $\mathbf{32.93\pm 0.00}$ $\mathbf{73.17\pm 0.00}$ $\mathbf{91.87\pm 0.00}$
$50$ ImageNet $7.36\pm 0.31$ $29.98\pm 1.73$ $67.58\pm 2.45$ $83.23\pm 0.34$
MoCoV3 $4.42\pm 0.00$ $39.43\pm 0.00$ $77.64\pm 0.00$ $92.68\pm 0.00$
SimCLRV2 $7.93\pm 0.00$ $37.40\pm 0.00$ $73.98\pm 0.00$ $91.87\pm 0.00$
DINO $7.62\pm 1.04$ $30.89\pm 1.22$ $62.81\pm 1.43$ $82.32\pm 3.05$
Our method $\mathbf{4.31\pm 0.01}$ $\mathbf{39.63\pm 0.61}$ $\mathbf{80.08\pm 0.41}$ $\mathbf{93.50\pm 0.41}$
Table 2: Comparing landmark detection test performance on Cephalometric dataset through our DDPM self-supervised pre-training and ImageNet, MoCoV3, SimCLRV2, DINO pre-training across varying numbers of labeled training images, highlighting best results in bold.
Number of Labeled Training Samples Pre-trained Models Cephalometric
MRE $\downarrow$ SDR(%) $\uparrow$
(mm) 2mm 2.5mm 3mm 4mm
$1$ ImageNet $48.69\pm 0.03$ $0.12\pm 0.01$ $0.16\pm 0.01$ $0.21\pm 0.02$ $0.41\pm 0.02$
MoCoV3 $24.39\pm 0.75$ $18.36\pm 0.35$ $23.92\pm 0.84$ $28.55\pm 0.93$ $35.89\pm 1.11$
SimCLRV2 $49.27\pm 9.00$ $4.14\pm 2.01$ $5.42\pm 2.56$ $6.73\pm 3.04$ $9.13\pm 4.10$
DINO $30.78\pm 0.73$ $7.64\pm 0.38$ $10.45\pm 0.50$ $13.57\pm 0.52$ $18.68\pm 0.64$
Our method $\mathbf{6.87\pm 0.00}$ $\mathbf{49.43\pm 0.00}$ $\mathbf{58.19\pm 0.00}$ $\mathbf{66.32\pm 0.00}$ $\mathbf{74.78\pm 0.00}$
$5$ ImageNet $46.84\pm 0.03$ $0.14\pm 0.02$ $0.22\pm 0.01$ $0.29\pm 0.00$ $0.52\pm 0.01$
MoCoV3 $7.53\pm 0.20$ $46.17\pm 0.41$ $55.26\pm 0.64$ $64.29\pm 0.53$ $75.48\pm 0.40$
SimCLRV2 $13.73\pm 0.16$ $33.69\pm 0.18$ $41.48\pm 0.35$ $49.31\pm 0.31$ $59.40\pm 0.10$
DINO $12.85\pm 0.22$ $35.88\pm 0.09$ $43.80\pm 0.27$ $50.65\pm 0.15$ $60.20\pm 0.32$
Our method $\mathbf{2.69\pm 0.00}$ $\mathbf{64.72\pm 0.00}$ $\mathbf{75.16\pm 0.00}$ $\mathbf{83.18\pm 0.00}$ $\mathbf{90.59\pm 0.00}$
$10$ ImageNet $27.44\pm 0.23$ $10.18\pm 0.03$ $13.33\pm 0.25$ $16.96\pm 0.20$ $22.68\pm 0.35$
MoCoV3 $5.32\pm 0.02$ $56.22\pm 0.29$ $66.91\pm 0.29$ $75.81\pm 0.13$ $85.25\pm 0.06$
SimCLRV2 $10.09\pm 0.07$ $43.67\pm 0.15$ $52.84\pm 0.09$ $61.19\pm 0.03$ $71.27\pm 0.20$
DINO $8.73\pm 0.08$ $48.19\pm 0.15$ $57.70\pm 0.27$ $65.90\pm 0.15$ $75.64\pm 0.00$
Our method $\mathbf{1.93\pm 0.00}$ $\mathbf{72.25\pm 0.00}$ $\mathbf{82.48\pm 0.00}$ $\mathbf{89.83\pm 0.00}$ $\mathbf{96.13\pm 0.00}$
$25$ ImageNet $4.18\pm 0.01$ $57.97\pm 0.37$ $69.03\pm 0.46$ $78.04\pm 0.20$ $87.11\pm 0.09$
MoCoV3 $2.22\pm 0.01$ $73.21\pm 0.20$ $82.69\pm 0.07$ $89.98\pm 0.25$ $95.79\pm 0.08$
SimCLRV2 $2.19\pm 0.01$ $69.35\pm 0.21$ $80.40\pm 0.04$ $88.36\pm 0.15$ $95.39\pm 0.17$
DINO $3.02\pm 0.06$ $68.22\pm 0.33$ $78.70\pm 0.27$ $86.81\pm 0.07$ $93.46\pm 0.09$
Our method $\mathbf{1.62\pm 0.03}$ $\mathbf{77.33\pm 1.12}$ $\mathbf{85.84\pm 1.49}$ $\mathbf{91.71\pm 1.20}$ $\mathbf{96.86\pm 0.63}$
$50$ ImageNet $1.60\pm 0.04$ $81.17\pm 0.42$ $89.20\pm 0.43$ $94.49\pm 0.32$ $98.29\pm 0.12$
MoCoV3 $1.63\pm 0.01$ $81.72\pm 0.04$ $89.72\pm 0.08$ $94.76\pm 0.02$ $98.28\pm 0.05$
SimCLRV2 $1.67\pm 0.02$ $79.55\pm 0.24$ $87.65\pm 0.27$ $93.67\pm 0.03$ $97.84\pm 0.01$
DINO $1.65\pm 0.01$ $80.89\pm 0.02$ $88.47\pm 0.48$ $93.75\pm 0.06$ $97.82\pm 0.07$
Our method $\mathbf{1.40\pm 0.01}$ $\mathbf{82.16\pm 1.16}$ $\mathbf{89.98\pm 0.62}$ $\mathbf{94.63\pm 0.54}$ $\mathbf{98.30\pm 0.09}$
Table 3: Comparing landmark detection test performance on the Hand dataset through our DDPM self-supervised pre-training and ImageNet, MoCoV3, SimCLRV2, and DINO pre-training across varying numbers of labeled training images, highlighting best results in bold.
Number of Labeled Training Samples Pre-trained Models Hand
MRE $\downarrow$ SDR(%) $\uparrow$
(mm) 2mm 4mm 10mm
$1$ ImageNet $79.32\pm 0.03$ $0.04\pm 0.00$ $0.14\pm 0.00$ $1.08\pm 0.02$
MoCoV3 $100.17\pm 0.10$ $0.42\pm 0.02$ $0.83\pm 0.02$ $1.60\pm 0.03$
SimCLRV2 $98.52\pm 0.09$ $0.06\pm 0.01$ $0.31\pm 0.00$ $1.02\pm 0.04$
DINO $97.31\pm 0.09$ $0.47\pm 0.04$ $1.27\pm 0.03$ $3.30\pm 0.18$
Our method $\mathbf{28.75\pm 0.00}$ $\mathbf{27.87\pm 0.00}$ $\mathbf{46.44\pm 0.00}$ $\mathbf{58.75\pm 0.00}$
$5$ ImageNet $86.12\pm 0.03$ $0.01\pm 0.01$ $0.09\pm 0.01$ $1.01\pm 0.01$
MoCoV3 $61.20\pm 0.39$ $9.82\pm 0.16$ $18.94\pm 0.29$ $32.11\pm 0.61$
SimCLRV2 $75.07\pm 0.48$ $5.99\pm 0.07$ $11.87\pm 0.14$ $20.35\pm 0.34$
DINO $86.03\pm 0.64$ $2.05\pm 0.31$ $5.79\pm 0.55$ $11.96\pm 0.84$
Our method $\mathbf{6.86\pm 0.00}$ $\mathbf{61.09\pm 0.00}$ $\mathbf{83.30\pm 0.00}$ $\mathbf{90.62\pm 0.00}$
$10$ ImageNet $85.87\pm 0.00$ $0.01\pm 0.00$ $0.15\pm 0.00$ $0.99\pm 0.00$
MoCoV3 $55.43\pm 0.29$ $13.92\pm 0.17$ $26.00\pm 0.15$ $40.21\pm 0.20$
SimCLRV2 $55.58\pm 0.75$ $16.72\pm 0.36$ $29.45\pm 0.45$ $42.18\pm 0.72$
DINO $60.58\pm 0.53$ $11.09\pm 0.14$ $23.00\pm 0.12$ $35.47\pm 0.32$
Our method $\mathbf{4.94\pm 0.11}$ $\mathbf{62.55\pm 0.55}$ $\mathbf{86.08\pm 0.25}$ $\mathbf{94.28\pm 0.17}$
$25$ ImageNet $24.96\pm 0.47$ $42.79\pm 0.60$ $57.93\pm 0.28$ $71.68\pm 0.18$
MoCoV3 $40.03\pm 0.12$ $27.19\pm 0.09$ $44.30\pm 0.27$ $56.55\pm 0.05$
SimCLRV2 $40.02\pm 0.11$ $29.69\pm 0.06$ $47.03\pm 0.00$ $59.73\pm 0.11$
DINO $49.27\pm 0.30$ $21.09\pm 0.06$ $36.12\pm 0.10$ $48.76\pm 0.20$
Our method $\mathbf{2.39\pm 0.00}$ $\mathbf{72.99\pm 0.00}$ $\mathbf{93.18\pm 0.00}$ $\mathbf{97.87\pm 0.00}$
$50$ ImageNet $10.79\pm 0.23$ $61.98\pm 0.36$ $78.51\pm 0.26$ $90.04\pm 0.28$
MoCoV3 $23.77\pm 0.34$ $47.44\pm 0.10$ $67.96\pm 0.14$ $76.18\pm 0.23$
SimCLRV2 $23.95\pm 0.18$ $49.28\pm 0.30$ $68.56\pm 0.45$ $77.03\pm 0.69$
DINO $34.95\pm 0.59$ $34.06\pm 0.97$ $52.32\pm 1.20$ $64.21\pm 0.80$
Our method $\mathbf{1.74\pm 0.03}$ $\mathbf{78.04\pm 0.94}$ $\mathbf{96.35\pm 0.51}$ $\mathbf{99.02\pm 0.04}$
Figure 4: Comparison of landmark detection performance between our pre-training method and other self-supervised and ImageNet pre-training approaches on Chest, Cephalometric, and Hand x-ray test sets, using only 10 labeled training samples.
Table 4: Landmark detection test performance of our method, pre-trained on the Hand dataset and fine-tuned on the Chest / Cephalometric datasets, across varying numbers of labeled training images.
Number of Labeled Training Samples Hand pre-training $\rightarrow$ Chest fine-tuning Hand pre-training $\rightarrow$ Cephalometric fine-tuning
MRE±$\pm$±STD $\downarrow$ SDR(%)$\pm$±STD $\uparrow$ MRE±$\pm$±STD $\downarrow$ SDR(%)$\pm$±STD $\uparrow$
(px) 3px 6px 9px (mm) 2mm 2.5mm 3mm 4mm
$1$ $23.79\pm 0.00$ $14.23\pm 0.00$ $35.37\pm 0.00$ $52.03\pm 0.00$ $9.56\pm 0.28$ $30.45\pm 1.73$ $38.60\pm 1.78$ $47.75\pm 2.00$ $60.21\pm 1.73$
$5$ $5.86\pm 0.00$ $32.93\pm 0.00$ $65.04\pm 0.00$ $82.93\pm 0.00$ $5.42\pm 0.00$ $55.07\pm 0.00$ $64.55\pm 0.00$ $72.44\pm 0.00$ $79.71\pm 0.00$
$10$ $5.08\pm 0.00$ $34.55\pm 0.00$ $70.73\pm 0.00$ $85.37\pm 0.00$ $2.78\pm 0.00$ $63.52\pm 0.30$ $74.20\pm 0.62$ $82.50\pm 0.35$ $89.97\pm 0.35$
$25$ $4.39\pm 0.00$ $38.62\pm 0.00$ $77.24\pm 0.00$ $91.46\pm 0.00$ $1.70\pm 0.00$ $74.03\pm 0.01$ $83.97\pm 0.05$ $90.95\pm 0.04$ $96.26\pm 0.03$
$50$ $3.99\pm 0.00$ $41.87\pm 0.00$ $82.52\pm 0.00$ $92.68\pm 0.00$ $1.50\pm 0.01$ $79.90\pm 0.02$ $87.85\pm 0.02$ $93.12\pm 0.00$ $97.37\pm 0.00$
5.3 Impact of different pre-training datasets
In this section, we explore whether a different in-domain dataset can still boost the performance in the potential situation where both the number of available images and annotations are limited. To this purpose, we select the best DDPM model pre-trained on the larger dataset (Hand), and we fine-tune it on the other two smaller datasets (Chest and Cephalometric), with the same protocol as the previous experiments (three runs and different numbers of annotated training images for the downstream task). Table 4 reports the obtained results. Our experiments reveal the efficacy of using a larger in-domain dataset (Hand) for pre-training in scenarios with limited target task data.
Using a model pre-trained on the Hand dataset for the Chest landmark detection task yields better results compared to a model pre-trained on the Chest dataset itself for most training samples quantity scenarios (5 to 50 labeled images). Even with just one labeled sample, the Hand pre-trained model achieves comparable performance to the Chest pre-trained model.
For the Cephalometric landmark detection task, the Hand pre-trained model does not outperform the Cephalometric pre-trained model, but it does achieve competitive results across all the selected number of training samples. The performance gap between the two pre-training approaches is relatively small, indicating that the Hand pre-training still transfers useful features to the Cephalometric task. Importantly, our approach consistently outperforms other pre-training methods (ImageNet, MoCoV3, SimCLRV2, DINO), regardless of the number of labeled training samples, for both tasks. This underscores the value of in-domain pre-training using our approach, even when the pre-training dataset differs from the target task. Our goal is to utilize images from the same dataset as the downstream task, thereby simulating real-world scenarios with limited data and label availability. Furthermore, we focused on label-efficient pre-training while minimizing computational resource usage, given that training diffusion models on extensive unlabeled datasets demands substantial resources and time. Even though different larger unlabeled datasets might enhance pre-training outcomes, our main point is that using a limited unannotated dataset can still surpass self-supervised and supervised state-of-the-art pre-training techniques like ImageNet (which has one million images), when images and annotations availability are constrained.
6 Conclusions
Despite the widespread application of deep neural networks for medical image analysis, data scarcity in images and annotations represents a fundamental challenge for this domain. In this research, we propose and evaluate the impact of a pre-training approach based on self-supervision with DDPM for label-efficient landmark detection tasks in x-ray images, benchmarking it against the typically adopted supervised ImageNet pre-training and self-supervised MoCoV3, SimCLRV2, and DINO. Specifically, given an x-ray image dataset, we leverage the training set in a self-supervised fashion for pre-training a DDPM based on a Unet, further fine-tuning it on the available annotated training images. Our results on three popular benchmark datasets show that the proposed method outperforms ImageNet supervised pre-training and MoCoV3, SimCLRV2, DINO self-supervised pre-training in the context of few annotations available for the landmark detection task. Moreover, such improvement seems to be preserved when pre-training on a different in-domain dataset and fine-tuning on the downstream one, simulating the interesting case where both annotations and images are limited for the downstream task. We believe that the proposed method showcases the potential of self-supervised diffusion-based pre-training for the landmark detection task, potentially opening new research directions to effectively address data scarcity challenges in this domain.

### v3

@@
 Abstract
-In the last few years, deep neural networks have been extensively applied in the medical domain for different tasks, ranging from image classification and segmentation to landmark detection.
-However, the application of these technologies in the medical domain is often hindered by data scarcity, both in terms of available annotations and images.
-This study introduces a new self-supervised pre-training protocol based on diffusion models for landmark detection in x-ray images.
-Our results show that the proposed self-supervised framework can provide accurate landmark detection with a minimal number of available annotated training images (up to $50$), outperforming ImageNet supervised pre-training and state-of-the-art self-supervised pre-trainings for three popular x-ray benchmark datasets.
-To our knowledge, this is the first exploration of diffusion models for self-supervised learning in landmark detection, which may offer a valuable pre-training approach in few-shot regimes, for mitigating data scarcity.
+Deep neural networks have been extensively applied in the medical domain for various tasks, including image classification, segmentation, and landmark detection.
+However, their application is often hindered by data scarcity, both in terms of available annotations and images.
+This study introduces a novel application of denoising diffusion probabilistic models (DDPMs) to the landmark detection task, specifically addressing the challenge of limited annotated data in x-ray imaging.
+Our key innovation lies in leveraging DDPMs for self-supervised pre-training in landmark detection, a previously unexplored approach in this domain.
+This method enables accurate landmark detection with minimal annotated training data (as few as $50$ images), surpassing both ImageNet supervised pre-training and traditional self-supervised techniques across three popular x-ray benchmark datasets.
+To our knowledge, this work represents the first application of diffusion models for self-supervised learning in landmark detection, which may offer a valuable pre-training approach in few-shot regimes, for mitigating data scarcity.
+To bolster further development and reproducibility, we provide open access to our code and pre-trained models for a variety of x-ray related applications: https://github.com/Malga-Vision/DiffusionXray-FewShot-LandmarkDetection
 1 Introduction
-Landmark detection is the task of identifying anatomical keypoints of interest in images [1].
-Detecting key points can directly assist in angle measurements [2], skeletal measurements [3], and surgical planning [4].
-In recent years, deep neural networks have become extensively utilized for the landmark detection task [5, 6], and many fully supervised methods have been proposed [7, 8, 9, 10, 11, 12, 13, 14, 15].
-However, the limited availability of annotated data and the reliance on medical experts for annotations present challenges in applying deep learning to this domain.
-Specifically, annotating landmarks is labor-intensive, costly, and demands expert precision [8, 7, 9], particularly for critical tasks such as surgical treatment planning.
+Landmark detection, the task of identifying anatomical keypoints in images [14], plays a crucial role in various medical applications, including angle measurements [21], skeletal assessments [19], and surgical planning [8].
+Although deep neural networks have become prevalent in this domain [29, 22], and many fully supervised methods have been proposed [43, 44, 33, 15, 36, 9, 39, 34, 16], their effectiveness is often limited by the scarcity of annotated data, particularly in medical imaging where expert annotations are costly and time consuming [44, 43, 33].
 Consequently, real-world datasets typically contain very few annotated images, requiring the design of label-efficient training solutions [42].
-A transfer learning framework, typically fine-tuning ImageNet pre-trained models on medical downstream tasks, has been a common strategy to tackle this challenge [17, 18, 19].
-Whether specific in-domain pre-training may benefit the model performance on the downstream task is still debated, and recent works provide controversial results depending on the specific task and domain [20, 9].
-Nevertheless, involving in-domain data may be an effective alternative to ImageNet pre-training, especially in a self-supervised fashion to overcome the limited availability of annotations.
-Recent advances in self-supervised learning (SSL) methods such as MoCoV3 [21], SimCLRV2 [22], and DINO [23] have shown significant potential.
-These approaches enable models to learn robust representations from unlabeled data, thereby reducing the dependency on large labeled training datasets.
-In this context, our main contribution is the proposal of a novel few-shot self-supervised pre-training approach, based on denoising diffusion probabilistic models (DDPMs), specifically focusing on landmark detection for x-ray images, widely used for anatomical assessment and diagnosis, in a realistic and typical scenario where the number of available annotations and images is minimal (up to $50$).
-To our knowledge, this work is the first to employ diffusion models for self-supervision in landmark detection.
-We compare the proposed method with state-of-the-art alternatives including ImageNet supervised pre-training, MoCoV3, SimCLRV2, and DINO self-supervised pre-training, in terms of performance and robustness at the varying of labeled training available samples.
-Our results show that our method outperforms ImageNet supervised pre-training and the other self-supervised methods for the landmark detection task.
+To address this challenge, researchers have explored transfer learning approaches, typically fine-tuning ImageNet pre-trained models on medical tasks [18, 1, 38].
+However, the efficacy of in-domain pre-training remains a subject of debate, with recent studies yielding conflicting results across different tasks and domains [31, 33].
+Self-supervised learning (SSL) methods such as MoCoV3 [6], SimCLRV2 [5], and DINO [4] have emerged as promising alternatives, enabling models to learn robust representations from unlabeled data.
+These approaches have shown potential to reduce the dependency on large labeled datasets, especially for medical imaging applications.
+In this context, our primary contribution lies in the novel application of DDPM to the task of landmark detection in x-ray images.
+While DDPMs have demonstrated remarkable success in image generation tasks [7, 24], their potential for self-supervised pre-training in medical image analysis, particularly landmark detection, remains largely unexplored.
+We propose a few-shot self-supervised pre-training approach specifically tailored for landmark detection in x-ray images, widely used for anatomical assessment and diagnosis, addressing the realistic and typical scenario where the available annotations and images are minimal (up to $50$).
+To our knowledge, this work represents the first application of diffusion models for self-supervision in landmark detection.
+We conduct a comprehensive comparison of our method against state-of-the-art alternatives, including YOLO [44], ImageNet supervised pre-training and self-supervised approaches like MoCoV3, SimCLRV2, and DINO.
+Our evaluation focuses on performance and robustness across varying quantities of labeled training samples.
+Our results demonstrate that our DDPM-based approach outperforms both ImageNet supervised pre-training and other self-supervised methods in the landmark detection task.
 Furthermore, we perform a set of experiments to provide insights on whether the usage of a different in-domain pre-training dataset may benefit the downstream task, in a few shot regimes.
+To facilitate further research in x-ray imaging, we have made our code and pre-trained models (including DDPM, MoCoV3, SimCLRV2, and DINO) publicly available.
+These resources, trained on the x-ray datasets used in this study, can be accessed at https://github.com/Malga-Vision/DiffusionXray-FewShot-LandmarkDetection.
+This work not only introduces a novel application of DDPMs for the landmark detection task but also contributes to the ongoing discussion on effective pre-training strategies for medical imaging tasks with limited annotations.
 The remainder of the paper is organized as follows: Section 2 reviews related studies, particularly focusing on self-supervised pre-training in the medical domain and landmark detection.
 Section 3 describes our proposed approach.
-Section 4 gives an overview of the benchmark datasets employed in this research, details the implementation specifics, and explains the evaluation metrics.
+Section 4 gives an overview of the benchmark datasets employed in this research and explains the evaluation metrics.
 Lastly, Section 5 showcases our experimental results and insights, drawing some conclusions in Section 6.
+Figure 1: Schematic representation of our DDPM self-supervised landmark detection pipeline.
 2 Related Works
-The problem of few-shot learning is particularly relevant in the medical application domain, as obtaining both images and high-quality, unbiased annotations is costly and time-consuming.
-Consequently, several works leverage self-supervised pre-training for label efficiency [24, 25, 26, 27].
-Zhu et al.[27] propose UOD (Universal One-shot Detection), a domain-adaptive framework for one-shot anatomical landmark detection across multiple medical imaging domains.
-UOD uses contrastive learning and a domain-adaptive transformer, while reducing annotation burden.
+Few-shot learning is particularly relevant in medical applications due to the high cost and time-intensive nature of obtaining high-quality, unbiased annotations [27].
+To address this challenge, researchers have explored various approaches, with transfer learning and self-supervised learning emerging as two prominent strategies.
+In landmark detection, existing approaches have primarily used transfer learning to address data scarcity [30, 44, 33].
+Tiulpin et al.
+[30] applied transfer learning from low-budget annotations for knee x-ray landmark localization in osteoarthritis stages.
+Di Via et al.
+[33] investigated the potential benefits of in-domain data over ImageNet pre-training for x-ray image landmark detection.
+Zhu et al.
+[44] proposed GU2Net, a universal model for anatomical landmark detection in medical images, addressing limitations in dataset-dependent methods.
+While transfer learning has shown promise, recent advancements in self-supervised learning have opened new avenues for tackling data scarcity in medical imaging tasks [32].
+Yao et al.
+[40] developed a two-stage framework called Cascade Comparing to Detect (CC2D) that uses self-supervised learning and pseudo-label training.
+Zhu et al.
+[42] introduced UOD, a domain-adaptive framework using contrastive learning and a domain-adaptive transformer.
+Miao et al.
+[23] proposed FM-OSD, using pre-trained visual foundation models for anatomical landmark detection introducing distance-aware similarity learning loss, and bidirectional matching strategy to enhance detection accuracy.
+In parallel with these developments, diffusion models have gained traction in various medical imaging tasks [17], although their potential for landmark detection remains largely unexplored.
+Pinaya et al.
+[26] developed a latent diffusion model for for 3D brain MRI generation.
 Rousseau et al.
-[24] introduce a pre-training method using diffusion models for dental radiography segmentation.
+[28] used diffusion models for dental radiography segmentation.
 Baranchuk et al.
-[25] use diffusion models to capture high-level semantic information for semantic segmentation.
-Brempong et al.
-[26] propose Decoder Denoising Pretraining (DDeP) to enhance label efficiency in semantic segmentation models by emphasizing the importance of pre-training both encoder and decoder components with limited labeled examples.
-In the specific context of landmark detection, the existing works have mainly tackled the data scarcity issue by leveraging transfer learning techniques [28, 8, 9].
-Tiulpin et al.
-[28] use transfer learning from low-budget annotations for knee x-ray landmark localization in osteoarthritis stages.
-Di Via et al.
-[9] study whether in-domain data provide benefits over ImageNet pre-training for landmark detection in x-ray images.
-Zhu et al.
-[8] propose GU2Net, a universal model for anatomical landmark detection in medical images, addressing limitations in specialized, dataset-dependent methods.
-Unlike the previously cited works, in this paper, we examine the impact of self-supervised pre-training with diffusion models on landmark detection for x-ray images, focusing on a realistic and typical scenario where the number of available annotations and images is limited.
-To our knowledge, this is the first application of diffusion models for SSL in this specific task.
-Figure 1: Schematic representation of our DDPM self-supervised landmark detection pipeline.
+[2] employed diffusion models for semantic segmentation.
+Wyatt et al.
+[37] proposed AnoDDPM for anomaly detection in medical images.
+Building upon these advancements in both self-supervised learning and diffusion models, our work distinguishes itself by examining the impact of self-supervised pre-training with diffusion models on landmark detection for x-ray images in a realistic scenario with limited annotations and images, a common challenge in medical imaging.
+To our knowledge, this study represents the first application of diffusion models for self-supervised learning in the specific task of landmark detection in x-ray images.
+This novel approach bridges the gap between the established potential of diffusion models in other medical imaging tasks and the pressing need for label-efficient methods in landmark detection.
 3 Approach
@@
 Subsequently, this model is fine-tuned on the labeled training set for the downstream task, represented by the landmark detection.
-Further details on these steps are discussed in the following section.
-Pre-training using the DDPM: Denoising Diffusion Probabilistic Models [29] (DDPMs) have emerged as a powerful class of generative models, excelling in the synthesis of high-fidelity samples in diverse fields, including medical imaging [30, 31].
-Their training methodology involves a forward diffusion process that gradually transforms data into noise and a reverse process that learns to reverse this transformation, thereby generating new data samples from noise.
-In this work, we leverage a DDPM Unet architecture as a predictive model for noise.
-During pre-training, the DDPM Unet receives batches of images alongside their respective diffusion time steps, and it is tasked with predicting the perturbation in the data.
-Real-world landmark detection datasets are typically small-scale, thus, we expect our DPPM Unet to be pre-trained with small-scale unannotated datasets, while still being able to provide rich and general features for the downstream task.
-This allows us to keep computational resources and time under a budget, while providing an effective few-shot learning approach for landmark detection.
-Fine-tuning on the downstream task: The pre-trained Unet model integrated into the DDPM undergoes fine-tuning for landmark detection in a few-shot regime, thus using a few available annotated data for the downstream task.
-This step requires only modifying the final classification layer to align with the number of landmarks to be predicted.
-Although the first stage adopts a self-supervised strategy (with no annotations), the second phase employs a supervised approach by providing the model with ground-truth heat maps for each landmark to be predicted.
+Further details on these steps are discussed below.
+DDPMs background.
+Denoising Diffusion Probabilistic Models (DDPMs) [12] have emerged as a powerful class of generative models, excelling in the synthesis of high-fidelity samples in diverse fields, including medical imaging [7, 24].
+DDPMs comprise two key processes: forward diffusion and reverse diffusion.
+1. Forward Diffusion Process:
+The forward diffusion process is described by:
+$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}\mathbf{I})$ (1)
+Here, $q(x_{t}|x_{t-1})$ represents the probability distribution of $x_{t}$ given $x_{t-1}$.
+$\mathcal{N}$ denotes a Gaussian distribution with mean $\sqrt{1-\beta_{t}}x_{t-1}$ and variance $\beta_{t}\mathbf{I}$.
+$\beta_{t}$ is a variance schedule that increases with $t$, controlling the amount of noise added at each step.
+$\mathbf{I}$ is the identity matrix.
+The process can be expressed as a single step:
+$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I})$ (2)
+where $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$.
+This equation allows direct sampling of $x_{t}$ given $x_{0}$, with $\sqrt{\bar{\alpha}_{t}}x_{0}$ as the mean and $(1-\bar{\alpha}_{t})\mathbf{I}$ as the variance.
+2. Reverse Diffusion Process:
+The reverse diffusion process is modeled as:
+$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$ (3)
+Here, $p_{\theta}(x_{t-1}|x_{t})$ is the probability of $x_{t-1}$ given $x_{t}$.
+$\mu_{\theta}(x_{t},t)$ and $\Sigma_{\theta}(x_{t},t)$ are the mean and covariance predicted by a neural network with parameters $\theta$.
+The mean of the reverse process is estimated by:
+$\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)\right)$ (4)
+where $\epsilon_{\theta}(x_{t},t)$ is the noise predicted by the neural network.
+The terms $\frac{1}{\sqrt{\alpha_{t}}}$ and $\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}$ are scaling factors that adjust for the noise level at time t.
+The training objective is:
+$\mathcal{L}_{\text{simple}}=\mathbb{E}_{t,x_{0},\epsilon}\left[\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\|^{2}\right]$ (5)
+This loss function trains the model to predict the noise $\epsilon$ added at each timestep.
+$\epsilon_{\theta}(\cdot)$ is the noise predicted by the model, and $\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ is the noisy input to the model at time t.
+The expectation $\mathbb{E}_{t,x_{0},\epsilon}$ is taken over random timesteps, images, and noise samples.
+Pre-training using the DDPM.
+In this work, we leverage a DDPM U-Net architecture as a pre-training strategy for anatomical landmark detection using unannotated medical images.
+This approach captures multi-scale anatomical representations, with DDPM’s denoising process extracting features at various levels and U-Net’s skip connections preserving both fine details and global information.
+During pre-training, the model learns to generate new images with the same pixel distribution as the originals, enabling it to learn important anatomical features.
+Given the typical scarcity of labeled data in real-world landmark detection, our DDPM U-Net is designed for effective pre-training on small-scale unannotated datasets while maintaining computational efficiency and minimizing training time.
+This approach yields rich and general features for the downstream task, offering an effective few-shot learning framework for landmark detection.
+By leveraging unannotated medical data, our model develops a comprehensive understanding of anatomical structures, creating a robust foundation that can adapt to specific landmark detection tasks with minimal additional training.
+Ground-truth heatmaps generation.
+We formulate landmark detection as a classification problem over all image pixels to generate output heatmaps.
+This approach enables simultaneous prediction of multiple landmarks, with one heatmap per landmark.
+Specifically, our model outputs $N$ heatmaps, where $N$ is the number of landmarks to predict.
+For heatmap generation, we adopt a Gaussian-based approach inspired by [44].
+This strategy involves applying a Gaussian filter ($\sigma=5$) centered at the ground-truth landmark location, followed by thresholding to obtain binary images.
+Formally:
+$H_{g}(i,j)=\begin{cases}1,&\text{if }\exp\left(-\frac{(i-x)^{2}+(j-y)^{2}}{2\sigma^{2}}\right)>\frac{1}{2}\max(G)\\ 0,&\text{otherwise}\end{cases}$
+Where $H_{g}(i,j)$ is the Gaussian heatmap pixel value at $(i,j)$, $(x,y)$ are the ground truth landmark coordinates, and $G$ is the Gaussian-filtered image before thresholding.
+Training on these heatmaps enables our model to learn landmark spatial relationships, predicting both general areas of interest and precise coordinates, and potentially enhance downstream landmark detection.
+Fine-tuning on the downstream task.
+The pre-trained DDPM UNet model is adapted for anatomical landmark detection, bridging the gap between generative modeling and precise anatomical localization.
+This adaptation, illustrated in Figure 1, involves a subtle yet powerful modification of the UNet architecture.
+While preserving the core structure that proved so effective in the DDPM, we alter the final convolutional layer to output multiple channels (each corresponding to a predicted heatmap for a specific anatomical landmark) and the timestep embedding is set to be null (as the temporal aspect is irrelevant for static landmark detection).
+This architectural shift transforms the model from a general image generator to a specialized landmark detector, capitalizing on learned hierarchical representations.
+Our fine-tuning strategy transitions from DDPM’s self-supervised paradigm to a supervised approach tailored for landmark detection.
+Training with ground-truth heatmaps allows the model to repurpose its learned features for precise spatial localization.
+Landmark coordinates are derived by calculating the centroid of each heatmap generated.
+This supervised fine-tuning, combined with pre-trained weight transfer (indicated by the blue dashed arrow), enables quick adaptation to the new task with minimal annotated data.
+The approach’s efficacy in few-shot learning contexts highlights its potential for medical imaging applications, where large annotated datasets are often scarce.
+By harnessing the power of diffusion models, traditionally used for generation tasks, we show their unexplored potential in discriminative tasks like landmark detection.
 4 Experimental Setup
-In this section, we describe the datasets used, the metrics employed, and the details of the implementation of our DDPM architectures and training procedures for the different pre-training methods discussed above.
-Afterwards, we show the results that have been achieved by evaluating the approach presented.
-4.1 Datasets Overview
-This section outlines the datasets employed in our study, including Chest x-rays, Cephalometric x-rays, and Hand x-rays.
-We describe the dataset composition, annotation protocols, and the division into training, validation, and testing sets to ensure clarity and reproducibility in our experiments.
-Chest dataset: This dataset comprises $279$ Chest x-ray images sourced from the China set, as documented in [32], and is accessible through the U.S.
-National Library of Medicine’s public repository.
-The original image dimensions are approximately $3000$x3000300030003000 pixels.
-Following the procedures provided in [8, 9], we split the dataset with the first $195$ images for training, the subsequent $34$ for validation and the final $50$ for testing.
-We evaluate our model’s performance by pixel distance using $6$ manually annotated landmarks per image as all the PNG dataset has no information on physical spacing.
-Cephalometric dataset: The Cephalometric dataset [33] is an open-source resource containing $400$ cephalometric x-ray images of individuals aged between $7$ and $76$ years.
-Each image has a resolution of $0.1$mmx0.10.10.10.1mm and dimensions of $2400$x1935193519351935 pixels.
-In our research, the first $150$ images are used for training purposes, while the remaining $250$ images are set aside for testing, following the methodology in [8].
-To ensure accurate anatomical landmark annotations, two orthodontists identified $19$ landmarks on each image, and the mean coordinates were calculated to address inter-observer variability.
-Hand dataset: This open dataset entails $909$ Hand x-ray images, each at an average size of $1563$x2169216921692169 pixels.
-Consistent with the approaches taken by [8, 9], we allocate the first $550$ images for training, the subsequent $59$ for validation, and the final $300$ images for testing.
-To enable a direct comparison with other approaches in terms of physical distances we take the estimate made by Payer et al.[34] that the width of the wrist is $50$mm.
-Then, physical distance is computed as $\frac{50}{|p-q|_{2}}$, where $p$ and $q$ are the two ends of the wrist, determined by first and fifth landmarks in the set of $37$ manually labelled by Payer et al.
-[34] as well.
-(a) Chest landmarks: 6
-(b) Cephalometric landmarks: 19
-(c) Hand landmarks: 37
-Figure 2: Annotated landmarks in x-ray images for Chest, Cephalometric, and Hand datasets.
+4.1 Datasets and Evaluation Metrics
+We evaluate our method on three public x-ray datasets standard in landmark detection research, following the dataset splits and evaluation protocols of [44, 33].
+The Chest x-ray dataset [3] comprises $279$ images (approximately $3000$×$3000$ pixels, resolution unknown) with $6$ landmarks, split into $195$ for training, $34$ for validation, and $50$ for testing.
+The Cephalometric x-ray dataset [35] consists of $400$ lateral cephalograms ($2400$×$1935$ pixels, $0.1$ mm resolution) with $19$ target landmarks, split into $130$ for training, $20$ for validation, and $250$ for testing.
+The Hand x-ray dataset [10] contains $909$ radiographs (average size $1563$×$2169$ pixels) with $37$ labeled landmarks, assuming a $50$ mm length between wrist endpoints [25], split into $550$ for training, $59$ for validation, and $300$ for testing.
+Our model is assessed using two standard metrics [27, 41]: the mean radial error (MRE) and the successful detection rate (SDR).
+MRE measures the Euclidean distance between predicted and ground truth landmarks, while SDR calculates the percentage of predictions within dataset-specific thresholds.
+For the Chest dataset, we report these metrics in pixels due to unavailable physical spacing information, whereas for Cephalometric and Hand datasets, we use millimeters, adhering to field standards.
 4.2 Implementation Details
-This section provides the technical specifics of the implementation and experimental setup.
-The experiments were conducted on a single NVIDIA A30 GPU with 24 GB of RAM, using Python 3.10 and PyTorch 2.1.0.
-Data pre-processing: Images are resized to $256$x256256256256 pixels and normalized to a $[0-1]$ range.
-Post-resizing, random augmentations are applied during training, including rotations (uniformly between $-2$°and $2$°), scaling (from $-0.02$ to $0.02$), and translations (from $-0.02$ to $0.02$).
-Pre-training stage: The DDPM is trained with a batch size of $4$, with gradient accumulation every $8$ batches.
+The experiments are conducted on a single NVIDIA A30 GPU with 24 GB RAM, using Python 3.10 and PyTorch 2.1.0.
+For data pre-processing, images are resized to $256$x256256256256 pixels, normalized to $[0-1]$ range, and augmented during training with random rotations ($-2$°and $2$°), scaling ($-0.02$ to $0.02$), and translations ($-0.02$ to $0.02$).
+Pre-training stage.
+The DDPM is trained with a batch size of $4$, with gradient accumulation every $8$ batches.
 Training involves $500$ diffusion steps with a linear noise schedule.
@@
 It incorporates attention mechanisms at a resolution of $32$, uses $4$ attention heads per channel, and includes $4$ residual blocks at each resolution level.
-Concerning the implementation of other SSL and ImageNet pre-training techniques, the sources used are [35] and [36].
-Fine-tuning stage: The fine-tuning process is carried out using the AdamW optimizer for $200$ epochs, with an early stopping criterion.
+Concerning the implementation of other SSL and ImageNet pre-training techniques, the sources are [11] and [13].
+Fine-tuning stage.
+The fine-tuning process is carried out using the AdamW optimizer for $200$ epochs, with an early stopping criterion.
 The learning rate is adjusted using a ReduceLROnPlateau scheduler based on the validation loss, starting at an initial value of $1e^{-5}$.
 The batch size is set at $2$, with a gradient accumulation of $8$.
-We use cross-entropy loss for our semantic landmark detection task to predict different classes.
-4.3 Evaluation Metrics
-In our research, we adopt the same evaluation metrics chosen by [8, 28, 9, 7], namely Mean Radial Error (MRE) and Success Detection Rate (SDR), which are widely used to assess the performance of landmark detection algorithms.
-The MRE measures the accuracy of predicted landmarks by calculating the average Euclidean distance to the corresponding ground truth landmarks. For $n$ landmarks, the MRE is defined as:
-$MRE=\frac{1}{n}\sum_{i=1}^{n}\sqrt{(x_{i}-\hat{x}_{i})^{2}+(y_{i}-\hat{y}_{i})^{2}}$
-where $(x_{i},y_{i})$ represent the ground truth landmark coordinates and $(\hat{x}_{i},\hat{y}_{i})$ represent the predicted coordinates.
-SDR, in contrast, evaluates robustness by indicating the percentage of predicted landmarks that fall within a certain threshold distance from the actual landmarks.
-For $n$ landmarks with a threshold $t$, the SDR is calculated as:
-$SDR=\frac{1}{n}\sum_{i=1}^{n}I(d_{i}\leq t)$
-where $d_{i}$ is the Euclidean distance between the ground truth and the predicted landmark, and $I$ is an indicator function that returns $1$ if $d_{i}\leq t$ and 00 otherwise.
-5 Results and Analysis Discussion
-In this section, we present the experiments conducted to evaluate our proposed methods and the corresponding results.
-We detail the tuning of DDPM pre-training iterations, assess the performance for landmark detection with few labeled data, and analyze the impact of different pre-training datasets on the downstream task.
+We use cross-entropy loss for gaussian heatmaps, and a negative log-likelihood loss for the contour-hugging ones [20].
+5 Experimental Results
 5.1 Tuning the DDPM pre-training iterations
-The number of training iterations during the pre-training of the DDPM is a fundamental hyperparameter of our proposed approach.
-To evaluate the robustness and select the number of training iterations, we perform an experiment running three times the entire pipeline, considering $4k$, $6k$, $8k$, and $10k$ training iterations for the DDPM pre-training step.
-Figure 3 visualizes the landmark detection results obtained in the validation sets for the Chest, Cephalometric, and Hand datasets, in terms of MRE.
-Our findings suggest that the pre-training is quite robust, except for the $4k$ iterations where the model is likely still learning crucial data features (as shown by the high standard deviation).
-Excessive training iterations beyond $8k$ may result in overfitting, and a slightly degrading performance on the downstream task.
-Considering these results, we decided to select the optimal number of training iterations for DDPM pre-training for subsequent experiments, utilizing $6k$ iterations for the Chest dataset, $8k$ iterations for the Cephalometric dataset, and $8k$ iterations for the Hand dataset.
+The number of training iterations during DDPM pre-training is a fundamental hyperparameter in our approach.
+To evaluate its impact and determine the optimal value, we conducted an experiment running the entire pipeline three times, considering $4k$, $6k$, $8k$, and $10k$ iterations for the DDPM pre-training step.
+Figure 2 illustrates the landmark detection results obtained on the validation sets for the Chest, Cephalometric, and Hand datasets in terms of MRE.
+Our findings indicate that the pre-training is quite robust, except for the $4k$ iterations where the model is likely still learning crucial data features (as shown by the high standard deviation).
+Excessive training iterations beyond $8k$ may result in overfitting, resulting in slightly degraded performance on the downstream task.
+Based on these results, we selected the optimal number of training iterations for DDPM pre-training in subsequent experiments, utilizing $6k$, $8k$, and $8k$ iterations for the Chest, Cephalometric, and Hand datasets respectively.
 Figure 2: Landmark detection performance on Chest, Cephalometric, and Hand validation sets across DDPM pre-training iterations.
 5.2 Downstream task performance evaluation
-In this section, we assess the effectiveness of our DDPM self-supervised pre-training method, benchmarking it with supervised ImageNet pre-training, and self-supervised state-of-the-art methods including MoCoV3, SimCLRV2, and DINO, across different numbers of labeled training samples (up to $50$) in the Chest, the Cephalometric, and the Hand datasets.
-Tables 1, 2, and 3 report the average test results for three independent runs on the three datasets (a bar chart version of the tables is provided in the Supplementary Material).
-With the exception of our approach, which implements a DDPM as described in the original paper [29], all reported results utilize DenseNet161 as the encoder backbone.
-We opt for this backbone due to its performance in the landmark detection task across all the datasets tested, based on a hold-out validation comparison with VGG19 and ResNeXt50_32x4D.
-Our approach outperforms both ImageNet and the explored alternative SSL approaches for all the tested datasets and quantities of training images in the downstream task.
-These results are qualitatively supported by Figure 4, providing a visual representation (i.e., the prediction heatmaps) of the methods’ performance with $10labels$, for each of the datasets included in our work.
-Performance gains are particularly pronounced in low-data regimes, crucial for medical imaging applications where labeled data is often scarce.
-In the Chest dataset (Table 1), with just one labeled sample, our method achieves an MRE of $14.99$ px compared to $143.67$ px for ImageNet and $64.69$ px for the best alternative (DINO).
-This represents a $89.6$% reduction in MRE versus ImageNet and a $76.8$% reduction versus DINO.
-The SDR at $9$px shows even more improvements, with our method reaching $64.63$%, far surpassing ImageNet (00%) and DINO ($23.44$%).
-Similar trends are observed in the Cephalometric dataset (Table 2).
-Using one labeled sample, our method attains an MRE of $6.87$ mm versus $48.69$ mm for ImageNet and $24.39$ mm for the best alternative (MoCoV3) - reductions of $85.9$% and $71.8$%, respectively.
-The SDR at the restrictive 2mm further highlights our method’s superiority, achieving $49.43$% compared to ImageNet’s $0.12$% and MoCoV3’s $18.36$%.
-The Hand dataset (Table 3) results reinforce these findings.
-With one labeled sample, our approach yields an MRE of $28.75$ mm versus $79.32$ mm for ImageNet and $97.31$ mm for the best alternative (DINO) - reductions of $63.8$% and $70.5$%, respectively.
-The SDR at 2mm shows our method at $27.87$%, significantly outperforming both ImageNet ($0.04$%) and DINO ($0.47$%).
-As the number of labeled samples increases beyond $10$, the performance gap narrows, though our approach still maintains an edge.
-For instance, in the Cephalometric dataset with $25$ labeled samples, our method’s MRE of $1.62$ mm is closer to, but still better than, ImageNet’s $4.18$ mm and SimCLRV2’s $2.19$ mm.
-This trend is consistent across all datasets, indicating that while our method provides substantial benefits in extremely low-data scenarios, for which it is conceived and proposed, it continues to offer improvements even as more labeled data becomes available.
-The obtained results highlight the efficacy of our DDPM self-supervised pre-training approach for landmark detection in few-shot learning scenarios, prevalent in medical imaging contexts where available high-quality annotated images are typically limited.
-Finally, to frame our results in the state-of-the-art, we can consider that the current best-performing method [9] for the landmark detection task of the whole Chest, Cephalometric, and Hand datasets, consists of fine-tuning an ImageNet pre-trained model on the target datasets.
+We assess the effectiveness of our DDPM self-supervised pre-training method by benchmarking it against supervised ImageNet pre-training and self-supervised state-of-the-art methods (MoCoV3, SimCLRV2, and DINO) across different numbers of labeled training samples (1, 5, 10, 25, and 50) in the Chest, Cephalometric, and Hand datasets.
+Additionally, we explore an ablation of our method using the contour-hugging approach [20] as an alternative framework for the downstream task.
+Tables 1, 2, and 3 report the average test results for three independent runs (a bar chart version of the tables is provided in the Supplementary Material).
+Except for our approach, which implements a DDPM as described in [12], all methods use DenseNet161 as the encoder backbone, chosen for its performance across all datasets based on hold-out validation comparisons with VGG19 and ResNeXt50_32x4D.
+Our approach consistently outperforms both ImageNet and alternative SSL approaches across all datasets and training image quantities.
+Figure 3 provides visual support, showing prediction heatmaps for each method with 10 labels across all datasets.
+Performance gains are particularly pronounced in low-data regimes.
+For instance, in the Chest dataset (Table 1) with one labeled sample, our method achieves an MRE of $14.99$px compared to $143.67$px for ImageNet and $64.69$px for DINO (the best alternative), representing reductions of $89.6$% and $76.8$% respectively.
+The SDR at $9$px shows even greater improvements, with our method reaching $64.63$%, far surpassing other approaches.
 Table 1: Comparing landmark detection test performance on the Chest dataset through our DDPM self-supervised pre-training and ImageNet, MoCoV3, SimCLRV2, DINO pre-training across varying numbers of labeled training images, highlighting best results in bold.
@@
 DINO $64.69\pm 0.91$ $3.66\pm 0.33$ $15.72\pm 0.38$ $23.44\pm 0.38$
-Our method $\mathbf{14.99\pm 0.00}$ $\mathbf{19.92\pm 0.00}$ $\mathbf{46.34\pm 0.00}$ $\mathbf{64.63\pm 0.00}$
+Our method with [20] $50.03\pm 0.00$ $\mathbf{21.14\pm 0.00}$ $43.90\pm 0.00$ $50.81\pm 0.00$
+Our method $\mathbf{14.99\pm 0.00}$ $19.92\pm 0.00$ $\mathbf{46.34\pm 0.00}$ $\mathbf{64.63\pm 0.00}$
 $5$ ImageNet $72.81\pm 2.71$ $2.44\pm 0.33$ $8.54\pm 1.00$ $14.09\pm 1.38$
@@
 DINO $24.48\pm 1.73$ $11.25\pm 1.16$ $32.65\pm 0.51$ $48.78\pm 1.00$
-Our method $\mathbf{6.04\pm 0.00}$ $\mathbf{32.52\pm 0.00}$ $\mathbf{65.04\pm 0.00}$ $\mathbf{86.99\pm 0.00}$
+Our method with [20] $8.14\pm 0.00$ $\mathbf{60.16\pm 0.00}$ $\mathbf{82.11\pm 0.00}$ $\mathbf{86.99\pm 0.00}$
+Our method $\mathbf{6.04\pm 0.00}$ $32.52\pm 0.00$ $65.04\pm 0.00$ $\mathbf{86.99\pm 0.00}$
 $10$ ImageNet $35.83\pm 1.57$ $7.11\pm 0.84$ $22.56\pm 1.06$ $37.30\pm 1.39$
@@
 DINO $14.96\pm 0.03$ $16.26\pm 0.41$ $39.64\pm 0.21$ $61.99\pm 0.61$
-Our method $\mathbf{5.33\pm 0.00}$ $\mathbf{33.74\pm 0.00}$ $\mathbf{67.89\pm 0.00}$ $\mathbf{86.59\pm 0.00}$
+Our method with [20] $\mathbf{4.47\pm 0.00}$ $\mathbf{60.98\pm 0.00}$ $\mathbf{88.62\pm 0.00}$ $\mathbf{93.50\pm 0.00}$
+Our method $5.33\pm 0.00$ $33.74\pm 0.00$ $67.89\pm 0.00$ $86.59\pm 0.00$
 $25$ ImageNet $11.38\pm 1.11$ $20.83\pm 2.00$ $49.29\pm 1.13$ $70.33\pm 0.29$
@@
 DINO $10.32\pm 0.35$ $24.59\pm 1.02$ $55.69\pm 0.81$ $75.81\pm 0.20$
-Our method $\mathbf{4.62\pm 0.00}$ $\mathbf{32.93\pm 0.00}$ $\mathbf{73.17\pm 0.00}$ $\mathbf{91.87\pm 0.00}$
+Our method with [20] $\mathbf{2.92\pm 0.00}$ $\mathbf{76.02\pm 0.00}$ $\mathbf{94.31\pm 0.00}$ $\mathbf{96.34\pm 0.00}$
+Our method $4.62\pm 0.00$ $32.93\pm 0.00$ $73.17\pm 0.00$ $91.87\pm 0.00$
 $50$ ImageNet $7.36\pm 0.31$ $29.98\pm 1.73$ $67.58\pm 2.45$ $83.23\pm 0.34$
@@
 DINO $7.62\pm 1.04$ $30.89\pm 1.22$ $62.81\pm 1.43$ $82.32\pm 3.05$
-Our method $\mathbf{4.31\pm 0.01}$ $\mathbf{39.63\pm 0.61}$ $\mathbf{80.08\pm 0.41}$ $\mathbf{93.50\pm 0.41}$
+Our method with [20] $\mathbf{2.69\pm 0.00}$ $\mathbf{76.83\pm 0.00}$ $\mathbf{93.50\pm 0.00}$ $\mathbf{97.56\pm 0.00}$
+Our method $4.31\pm 0.01$ $39.63\pm 0.61$ $80.08\pm 0.41$ $93.50\pm 0.41$
+Similar trends are observed in the Cephalometric (Table 2) and Hand datasets (Table 3).
+With one labeled sample in the Cephalometric dataset, our method attains an MRE of $15.71$mm versus $86.71$mm for ImageNet and $43.49$mm for MoCoV3, with corresponding improvements in SDR.
+In the Hand dataset, our approach yields an MRE of $28.75$mm versus $79.32$mm for ImageNet and $97.31$mm for DINO, again with SDR improvements.
+As the number of labeled samples increases, the performance gap narrows, though our approach maintains an edge.
+For example, in the Cephalometric dataset with $25$ labeled samples, our method’s MRE of $2.84$mm still outperforms ImageNet’s $7.67$mm and SimCLRV2’s $3.99$mm.
+This trend is consistent across all datasets, indicating that our method provides substantial benefits in extremely low-data scenarios while continuing to offer improvements as more labeled data becomes available.
+These results highlight our DDPM self-supervised pre-training’s effectiveness for landmark detection in few-shot learning scenarios, common in medical imaging where high-quality annotated images are limited.
 Table 2: Comparing landmark detection test performance on Cephalometric dataset through our DDPM self-supervised pre-training and ImageNet, MoCoV3, SimCLRV2, DINO pre-training across varying numbers of labeled training images, highlighting best results in bold.
@@
 (mm) 2mm 2.5mm 3mm 4mm
-$1$ ImageNet $48.69\pm 0.03$ $0.12\pm 0.01$ $0.16\pm 0.01$ $0.21\pm 0.02$ $0.41\pm 0.02$
-MoCoV3 $24.39\pm 0.75$ $18.36\pm 0.35$ $23.92\pm 0.84$ $28.55\pm 0.93$ $35.89\pm 1.11$
-SimCLRV2 $49.27\pm 9.00$ $4.14\pm 2.01$ $5.42\pm 2.56$ $6.73\pm 3.04$ $9.13\pm 4.10$
-DINO $30.78\pm 0.73$ $7.64\pm 0.38$ $10.45\pm 0.50$ $13.57\pm 0.52$ $18.68\pm 0.64$
-Our method $\mathbf{6.87\pm 0.00}$ $\mathbf{49.43\pm 0.00}$ $\mathbf{58.19\pm 0.00}$ $\mathbf{66.32\pm 0.00}$ $\mathbf{74.78\pm 0.00}$
-$5$ ImageNet $46.84\pm 0.03$ $0.14\pm 0.02$ $0.22\pm 0.01$ $0.29\pm 0.00$ $0.52\pm 0.01$
-MoCoV3 $7.53\pm 0.20$ $46.17\pm 0.41$ $55.26\pm 0.64$ $64.29\pm 0.53$ $75.48\pm 0.40$
-SimCLRV2 $13.73\pm 0.16$ $33.69\pm 0.18$ $41.48\pm 0.35$ $49.31\pm 0.31$ $59.40\pm 0.10$
-DINO $12.85\pm 0.22$ $35.88\pm 0.09$ $43.80\pm 0.27$ $50.65\pm 0.15$ $60.20\pm 0.32$
-Our method $\mathbf{2.69\pm 0.00}$ $\mathbf{64.72\pm 0.00}$ $\mathbf{75.16\pm 0.00}$ $\mathbf{83.18\pm 0.00}$ $\mathbf{90.59\pm 0.00}$
-$10$ ImageNet $27.44\pm 0.23$ $10.18\pm 0.03$ $13.33\pm 0.25$ $16.96\pm 0.20$ $22.68\pm 0.35$
-MoCoV3 $5.32\pm 0.02$ $56.22\pm 0.29$ $66.91\pm 0.29$ $75.81\pm 0.13$ $85.25\pm 0.06$
-SimCLRV2 $10.09\pm 0.07$ $43.67\pm 0.15$ $52.84\pm 0.09$ $61.19\pm 0.03$ $71.27\pm 0.20$
-DINO $8.73\pm 0.08$ $48.19\pm 0.15$ $57.70\pm 0.27$ $65.90\pm 0.15$ $75.64\pm 0.00$
-Our method $\mathbf{1.93\pm 0.00}$ $\mathbf{72.25\pm 0.00}$ $\mathbf{82.48\pm 0.00}$ $\mathbf{89.83\pm 0.00}$ $\mathbf{96.13\pm 0.00}$
-$25$ ImageNet $4.18\pm 0.01$ $57.97\pm 0.37$ $69.03\pm 0.46$ $78.04\pm 0.20$ $87.11\pm 0.09$
-MoCoV3 $2.22\pm 0.01$ $73.21\pm 0.20$ $82.69\pm 0.07$ $89.98\pm 0.25$ $95.79\pm 0.08$
-SimCLRV2 $2.19\pm 0.01$ $69.35\pm 0.21$ $80.40\pm 0.04$ $88.36\pm 0.15$ $95.39\pm 0.17$
-DINO $3.02\pm 0.06$ $68.22\pm 0.33$ $78.70\pm 0.27$ $86.81\pm 0.07$ $93.46\pm 0.09$
-Our method $\mathbf{1.62\pm 0.03}$ $\mathbf{77.33\pm 1.12}$ $\mathbf{85.84\pm 1.49}$ $\mathbf{91.71\pm 1.20}$ $\mathbf{96.86\pm 0.63}$
-$50$ ImageNet $1.60\pm 0.04$ $81.17\pm 0.42$ $89.20\pm 0.43$ $94.49\pm 0.32$ $98.29\pm 0.12$
-MoCoV3 $1.63\pm 0.01$ $81.72\pm 0.04$ $89.72\pm 0.08$ $94.76\pm 0.02$ $98.28\pm 0.05$
-SimCLRV2 $1.67\pm 0.02$ $79.55\pm 0.24$ $87.65\pm 0.27$ $93.67\pm 0.03$ $97.84\pm 0.01$
-DINO $1.65\pm 0.01$ $80.89\pm 0.02$ $88.47\pm 0.48$ $93.75\pm 0.06$ $97.82\pm 0.07$
-Our method $\mathbf{1.40\pm 0.01}$ $\mathbf{82.16\pm 1.16}$ $\mathbf{89.98\pm 0.62}$ $\mathbf{94.63\pm 0.54}$ $\mathbf{98.30\pm 0.09}$
+$1$ ImageNet $86.71\pm 0.02$ $0.02\pm 0.00$ $0.06\pm 0.00$ $0.06\pm 0.00$ $0.15\pm 0.00$
+MoCoV3 $43.49\pm 1.21$ $6.79\pm 0.10$ $10.83\pm 0.25$ $13.67\pm 0.30$ $19.99\pm 0.30$
+SimCLRV2 $100.68\pm 12.18$ $0.82\pm 0.42$ $1.27\pm 0.68$ $1.63\pm 0.79$ $2.51\pm 1.19$
+DINO $58.67\pm 1.22$ $2.47\pm 0.14$ $4.42\pm 0.07$ $5.52\pm 0.07$ $8.60\pm 0.30$
+Our method with [20] $20.13\pm 0.00$ $\mathbf{18.06\pm 0.00}$ $25.14\pm 0.00$ $31.94\pm 0.00$ $41.37\pm 0.00$
+Our method $\mathbf{15.71\pm 0.00}$ $17.31\pm 0.00$ $\mathbf{27.14\pm 0.00}$ $\mathbf{33.24\pm 0.00}$ $\mathbf{45.14\pm 0.00}$
+$5$ ImageNet $84.40\pm 0.02$ $0.06\pm 0.00$ $0.11\pm 0.00$ $0.15\pm 0.00$ $0.17\pm 0.00$
+MoCoV3 $13.44\pm 0.12$ $18.62\pm 0.19$ $29.56\pm 0.18$ $35.99\pm 0.02$ $49.38\pm 0.16$
+SimCLRV2 $24.59\pm 0.12$ $13.06\pm 0.31$ $21.24\pm 0.25$ $26.44\pm 0.23$ $36.52\pm 0.37$
+DINO $23.83\pm 0.21$ $14.00\pm 0.26$ $23.01\pm 0.54$ $28.42\pm 0.52$ $39.13\pm 0.38$
+Our method with [20] $11.53\pm 0.00$ $\mathbf{45.05\pm 0.00}$ $\mathbf{54.00\pm 0.00}$ $\mathbf{61.03\pm 0.00}$ $\mathbf{70.55\pm 0.00}$
+Our method $\mathbf{8.30\pm 0.00}$ $27.18\pm 0.00$ $42.65\pm 0.00$ $51.01\pm 0.00$ $66.34\pm 0.00$
+$10$ ImageNet $50.09\pm 0.44$ $3.18\pm 0.16$ $5.92\pm 0.25$ $7.49\pm 0.26$ $11.19\pm 0.42$
+MoCoV3 $9.30\pm 0.03$ $23.84\pm 0.12$ $37.20\pm 0.13$ $45.34\pm 0.17$ $60.18\pm 0.13$
+SimCLRV2 $18.54\pm 0.09$ $17.80\pm 0.29$ $28.16\pm 0.04$ $34.65\pm 0.22$ $46.74\pm 0.23$
+DINO $15.64\pm 0.26$ $18.86\pm 0.12$ $30.89\pm 0.10$ $38.22\pm 0.34$ $51.83\pm 0.45$
+Our method with [20] $5.21\pm 0.00$ $\mathbf{52.11\pm 0.00}$ $\mathbf{62.97\pm 0.00}$ $\mathbf{71.43\pm 0.00}$ $\mathbf{81.68\pm 0.00}$
+Our method $\mathbf{3.52\pm 0.00}$ $32.06\pm 0.00$ $48.59\pm 0.00$ $58.97\pm 0.00$ $74.69\pm 0.00$
+$25$ ImageNet $7.67\pm 0.39$ $24.25\pm 0.10$ $37.83\pm 0.13$ $45.58\pm 0.26$ $62.56\pm 0.35$
+MoCoV3 $4.04\pm 0.03$ $34.81\pm 0.18$ $51.26\pm 0.14$ $60.80\pm 0.03$ $76.78\pm 0.10$
+SimCLRV2 $3.99\pm 0.04$ $30.63\pm 0.23$ $47.04\pm 0.18$ $56.20\pm 0.13$ $73.29\pm 0.21$
+DINO $5.38\pm 0.19$ $30.00\pm 0.20$ $45.98\pm 0.04$ $55.36\pm 0.06$ $72.30\pm 0.19$
+Our method with [20] $\mathbf{2.72\pm 0.00}$ $\mathbf{65.47\pm 0.00}$ $\mathbf{75.77\pm 0.00}$ $\mathbf{82.55\pm 0.00}$ $\mathbf{90.61\pm 0.00}$
+Our method $2.84\pm 0.00$ $39.84\pm 0.01$ $57.61\pm 0.01$ $67.17\pm 0.01$ $81.01\pm 0.00$
+$50$ ImageNet $2.82\pm 0.09$ $42.93\pm 2.66$ $60.74\pm 3.48$ $69.69\pm 3.28$ $83.87\pm 2.24$
+MoCoV3 $2.98\pm 0.01$ $43.68\pm 0.32$ $61.84\pm 0.13$ $70.48\pm 0.03$ $84.81\pm 0.19$
+SimCLRV2 $3.17\pm 0.04$ $39.94\pm 0.19$ $58.70\pm 0.08$ $67.66\pm 0.05$ $81.85\pm 0.10$
+DINO $2.98\pm 0.02$ $40.33\pm 0.38$ $59.06\pm 0.39$ $68.99\pm 0.42$ $83.78\pm 0.64$
+Our method with [20] $\mathbf{2.13\pm 0.00}$ $\mathbf{73.60\pm 0.00}$ $\mathbf{81.22\pm 0.00}$ $\mathbf{86.78\pm 0.00}$ $\mathbf{92.70\pm 0.00}$
+Our method $2.50\pm 0.00$ $44.91\pm 0.05$ $62.84\pm 0.02$ $72.82\pm 0.00$ $85.81\pm 0.02$
 Table 3: Comparing landmark detection test performance on the Hand dataset through our DDPM self-supervised pre-training and ImageNet, MoCoV3, SimCLRV2, and DINO pre-training across varying numbers of labeled training images, highlighting best results in bold.
@@
 DINO $97.31\pm 0.09$ $0.47\pm 0.04$ $1.27\pm 0.03$ $3.30\pm 0.18$
+Our method with [20] $47.76\pm 0.00$ $8.70\pm 0.00$ $12.76\pm 0.00$ $16.59\pm 0.00$
 Our method $\mathbf{28.75\pm 0.00}$ $\mathbf{27.87\pm 0.00}$ $\mathbf{46.44\pm 0.00}$ $\mathbf{58.75\pm 0.00}$
@@
 DINO $86.03\pm 0.64$ $2.05\pm 0.31$ $5.79\pm 0.55$ $11.96\pm 0.84$
+Our method with [20] $37.53\pm 0.00$ $21.49\pm 0.00$ $28.90\pm 0.00$ $34.17\pm 0.00$
 Our method $\mathbf{6.86\pm 0.00}$ $\mathbf{61.09\pm 0.00}$ $\mathbf{83.30\pm 0.00}$ $\mathbf{90.62\pm 0.00}$
@@
 DINO $60.58\pm 0.53$ $11.09\pm 0.14$ $23.00\pm 0.12$ $35.47\pm 0.32$
+Our method with [20] $25.47\pm 0.00$ $35.67\pm 0.00$ $45.34\pm 0.00$ $51.01\pm 0.00$
 Our method $\mathbf{4.94\pm 0.11}$ $\mathbf{62.55\pm 0.55}$ $\mathbf{86.08\pm 0.25}$ $\mathbf{94.28\pm 0.17}$
@@
 DINO $49.27\pm 0.30$ $21.09\pm 0.06$ $36.12\pm 0.10$ $48.76\pm 0.20$
+Our method with [20] $11.99\pm 0.00$ $62.21\pm 0.00$ $77.03\pm 0.00$ $81.39\pm 0.00$
 Our method $\mathbf{2.39\pm 0.00}$ $\mathbf{72.99\pm 0.00}$ $\mathbf{93.18\pm 0.00}$ $\mathbf{97.87\pm 0.00}$
@@
 DINO $34.95\pm 0.59$ $34.06\pm 0.97$ $52.32\pm 1.20$ $64.21\pm 0.80$
-Our method $\mathbf{1.74\pm 0.03}$ $\mathbf{78.04\pm 0.94}$ $\mathbf{96.35\pm 0.51}$ $\mathbf{99.02\pm 0.04}$
-Figure 4: Comparison of landmark detection performance between our pre-training method and other self-supervised and ImageNet pre-training approaches on Chest, Cephalometric, and Hand x-ray test sets, using only 10 labeled training samples.
-Table 4: Landmark detection test performance of our method, pre-trained on the Hand dataset and fine-tuned on the Chest / Cephalometric datasets, across varying numbers of labeled training images.
+Our method with [20] $7.16\pm 0.00$ $\mathbf{79.38\pm 0.00}$ $88.36\pm 0.00$ $90.19\pm 0.00$
+Our method $\mathbf{1.74\pm 0.03}$ $78.04\pm 0.94$ $\mathbf{96.35\pm 0.51}$ $\mathbf{99.02\pm 0.04}$
+Figure 3: Comparison of landmark detection performance between our DDPM pre-training method and alternative self-supervised and ImageNet pre-training approaches on Chest, Cephalometric, and Hand x-ray test sets, using only 10 labeled training samples.
+Table 4: Comparison of our method vs.
+state-of-the-art YOLO framework using 1, 5, and 10 labeled samples.
++ indicates mixed dataset training.
+Both methods utilizes input images resized at $256$x256256256256 pixels.
+Our method performs competitively using only 1 labeled sample.
+Methods Number of Labeled Training Samples Chest Cephalometric Hand
+MRE $\downarrow$ SDR(%) $\uparrow$ MRE $\downarrow$ SDR(%) $\uparrow$ MRE $\downarrow$ SDR(%) $\uparrow$
+(px) 3px 6px 9px (mm) 2mm 2.5mm 3mm 4mm (mm) 2mm 4mm 10mm
+YOLO+ 10 22.55 32.11 63.82 76.42 15.05 18.23 26.50 39.62 53.49 52.43 11.59 19.52 24.14
+YOLO+ 5 23.13 23.57 50.40 67.07 19.38 11.53 17.07 25.95 35.53 64.24 15.15 24.91 31.79
+YOLO+ 1 84.37 4.07 13.41 19.51 53.80 7.35 11.22 17.26 23.39 61.36 8.26 17.16 31.47
+Our method 1 14.99 19.92 46.34 64.63 15.71 17.31 27.14 33.24 45.14 28.75 27.87 46.44 58.75
+Table 5: Landmark detection test performance of our Gaussian-heatmaps method, pre-trained on the Hand dataset and fine-tuned on the Chest / Cephalometric datasets, across varying numbers of labeled training images.
 Number of Labeled Training Samples Hand pre-training $\rightarrow$ Chest fine-tuning Hand pre-training $\rightarrow$ Cephalometric fine-tuning
@@
 $50$ $3.99\pm 0.00$ $41.87\pm 0.00$ $82.52\pm 0.00$ $92.68\pm 0.00$ $1.50\pm 0.01$ $79.90\pm 0.02$ $87.85\pm 0.02$ $93.12\pm 0.00$ $97.37\pm 0.00$
+Finally, to contextualize our results, we compare our method to YOLO [44], a state-of-the-art universal anatomical landmark detection model.
+YOLO utilizes mixed dataset training and combines a universal U-Net for local feature learning with a global network for landmark disambiguation.
+As Table 4 shows, our method performs comparably or better than YOLO variants across all three datasets.
+Despite YOLO’s mixed dataset training (denoted by ’$+$’), our approach achieves competitive results with just one labeled sample, even with $5$ and $10$ samples.
+Ablation study on the downstream task setting.
+We perform an ablation study comparing the performance of our downstream task framework, with the one introduced in [20] (second-last entry in our results Tables), mainly differing in the usage of a single annotated pixel to generate unrestricted contour-hugging heatmaps and a negative log-likelihood loss.
+For the Hand dataset, which had more extensive pre-training data, our approach consistently outperforms the evaluated alternative.
+However, the contour-hugging method generally shows better SDR performance on the Chest and the Cephalometric datasets.
+Regarding MRE, our original framework performs better up to $10$ training samples, showing lower performance with more annotated data.
+These findings suggest that the effectiveness of each downstream setting depends on pre-training data volume and quality, as well as dataset-specific characteristics.
 5.3 Impact of different pre-training datasets
-In this section, we explore whether a different in-domain dataset can still boost the performance in the potential situation where both the number of available images and annotations are limited.
-To this purpose, we select the best DDPM model pre-trained on the larger dataset (Hand), and we fine-tune it on the other two smaller datasets (Chest and Cephalometric), with the same protocol as the previous experiments (three runs and different numbers of annotated training images for the downstream task).
-Table 4 reports the obtained results.
-Our experiments reveal the efficacy of using a larger in-domain dataset (Hand) for pre-training in scenarios with limited target task data.
-Using a model pre-trained on the Hand dataset for the Chest landmark detection task yields better results compared to a model pre-trained on the Chest dataset itself for most training samples quantity scenarios (5 to 50 labeled images).
-Even with just one labeled sample, the Hand pre-trained model achieves comparable performance to the Chest pre-trained model.
-For the Cephalometric landmark detection task, the Hand pre-trained model does not outperform the Cephalometric pre-trained model, but it does achieve competitive results across all the selected number of training samples.
-The performance gap between the two pre-training approaches is relatively small, indicating that the Hand pre-training still transfers useful features to the Cephalometric task.
-Importantly, our approach consistently outperforms other pre-training methods (ImageNet, MoCoV3, SimCLRV2, DINO), regardless of the number of labeled training samples, for both tasks.
+In this section, we investigate whether a different in-domain dataset can enhance performance when both available images and annotations are limited.
+We select the best DDPM model pre-trained on the larger Hand dataset and fine-tune it on the smaller Chest and Cephalometric datasets, following the same protocol as previous experiments with Gaussian heatmaps.
+Table 5 reports the results.
+Our experiments show the effectiveness of using a larger in-domain dataset (Hand) for pre-training in scenarios with limited target task data.
+For the Chest landmark detection task, the Hand pre-trained model outperforms the Chest pre-trained model in most scenarios (5 to 50 labeled images).
+Even with just one labeled sample, it achieves comparable performance to the Chest pre-trained model.
+For the Cephalometric landmark detection task, while the Hand pre-trained model doesn’t surpass the Cephalometric pre-trained model, it achieves competitive results across all training sample quantities.
+The small performance gap indicates that Hand pre-training still transfers useful features to the Cephalometric task.
+Notably, our approach consistently outperforms other pre-training methods (ImageNet, MoCoV3, SimCLRV2, DINO) and frameworks (YOLO) for both tasks, regardless of the number of labeled training samples.
 This underscores the value of in-domain pre-training using our approach, even when the pre-training dataset differs from the target task.
-Our goal is to utilize images from the same dataset as the downstream task, thereby simulating real-world scenarios with limited data and label availability.
-Furthermore, we focused on label-efficient pre-training while minimizing computational resource usage, given that training diffusion models on extensive unlabeled datasets demands substantial resources and time.
-Even though different larger unlabeled datasets might enhance pre-training outcomes, our main point is that using a limited unannotated dataset can still surpass self-supervised and supervised state-of-the-art pre-training techniques like ImageNet (which has one million images), when images and annotations availability are constrained.
+Our goal is to simulate real-world scenarios with limited data and label availability by utilizing images from the same dataset as the downstream task.
+We focused on label-efficient pre-training while minimizing computational resource usage, given the substantial resources and time required for training diffusion models on extensive unlabeled datasets.
+While larger unlabeled datasets might enhance pre-training outcomes, our key finding is that using a limited unannotated dataset can still surpass self-supervised and supervised state-of-the-art pre-training techniques like ImageNet (which has one million images) when image and annotation availability is constrained.
 6 Conclusions
-Despite the widespread application of deep neural networks for medical image analysis, data scarcity in images and annotations represents a fundamental challenge for this domain.
-In this research, we propose and evaluate the impact of a pre-training approach based on self-supervision with DDPM for label-efficient landmark detection tasks in x-ray images, benchmarking it against the typically adopted supervised ImageNet pre-training and self-supervised MoCoV3, SimCLRV2, and DINO.
-Specifically, given an x-ray image dataset, we leverage the training set in a self-supervised fashion for pre-training a DDPM based on a Unet, further fine-tuning it on the available annotated training images.
-Our results on three popular benchmark datasets show that the proposed method outperforms ImageNet supervised pre-training and MoCoV3, SimCLRV2, DINO self-supervised pre-training in the context of few annotations available for the landmark detection task.
-Moreover, such improvement seems to be preserved when pre-training on a different in-domain dataset and fine-tuning on the downstream one, simulating the interesting case where both annotations and images are limited for the downstream task.
-We believe that the proposed method showcases the potential of self-supervised diffusion-based pre-training for the landmark detection task, potentially opening new research directions to effectively address data scarcity challenges in this domain.
+Our research introduces a novel application of DDPMs for anatomical landmark detection in X-ray images, addressing the persistent challenge of limited annotated data in medical image analysis.
+This work bridges the gap between the established potential of diffusion models in medical imaging and the critical need for label-efficient methods in landmark localization.
+We demonstrate that our DDPM-based self-supervised pre-training method consistently outperforms supervised ImageNet pre-training, traditional state-of-the-art self-supervised approaches (MoCoV3, SimCLRV2, DINO), and the YOLO framework in few-shot learning scenarios across Chest, Cephalometric, and Hand X-ray datasets.
+The performance gains are particularly significant in low-data regimes, with substantial MRE reductions and SDR improvements.
+Notably, our method maintains its effectiveness when pre-trained on one in-domain dataset and fine-tuned on smaller, distinct datasets, simulating real-world scenarios with limited data availability.
+This robustness suggests broad applicability across various medical imaging tasks and datasets.
+Our work not only addresses the critical issue of data scarcity but also contributes to the ongoing discussion on effective pre-training strategies in medical image analysis.
+By showing the efficacy of DDPMs in this context, we aims to open new avenues for research into efficient and robust landmark detection methods, potentially accelerating progress in automated diagnosis systems.